(a) Assuming Doll uses traditional cost accounting procedures:

1. How much cost was charged to Work-in-Process during the month?
$____

2. How much cost was charged to cost of goods sold during the month?
$____

(b) Assuming Doll is a lean production company and uses backflush costing method:

1. How much cost was charged to Work-in-Process during the month?
$_____

2. How much cost was charged to cost of goods sold during the month?
$______

Loberly manufactures laptops under its own brand, but acquires all the components from outside vendors. No computers are assembled until the order is received online from customers, so there is no finished goods inventory. When an order is received, the bill of materials required to fill the order is prepared automatically and sent electronically to the various vendors. All components are received from vendors within three days and the completed order is shipped to the customer immediately when completed, usually on the same day the components are received from vendors. The number of units in process at the end of any day is negligible.

The following data are provided for the most recent month of operations:

Actual components costs incurred $893,000
Actual conversion costs incurred 205,000
Units in process, beginning of month -0-
Units started in process during the month 8,000
Units in process, end of month -0-
Answer #1

PART A

1. cost charged to Work-in-Process during the month = (893000+205000) = 1098000

2.cost charged to cost of goods sold during the month = (893000+205000) = 1098000

PART B

1. cost charged to Work-in-Process during the month = 0 (because in  backflush costing method, nothing is charge to work in process)

2.cost charged to cost of goods sold during the month = (893000+205000) = 1098000